OpenStudy (anonymous):

The equation s = p - 0.15p represents the sale price (s) of an item with an original price (p), after a 15% discount. Make a table to find the discount prices for items with original prices of $12, $16, $20, and $25. -- Can someone please help? ;_; I've been stuck on this problem for a while now. :(

OpenStudy (anonymous):

okay the discounted price = original price - 0.15 *(original price ) = (1-0.15)* original price = 0.85* original price So for discounted price of $12, the asnwer will be 0.85*12 Similarly for the others also..

OpenStudy (anonymous):

Thank you! :) I'm homeschooled so I don't get things like this explained me, again thank you!! ^^;

OpenStudy (anonymous):

No Probs! :)